OpenStudy (anonymous):

Please help with this one too... solve Ed and Rick leave the same point at the same time. Ed is bicycling west at a rate of 2 mi/h faster than Rick, who is traveling south. After one hour, they are 10 miles apart. How fast is each person traveling?

OpenStudy (anonymous):

Can somebody help me???

OpenStudy (anonymous):

S1² + S2² = 100 mi (x* t)² + ( y *t)² = 100 x² + ( x+2 )² = 100 -> 2 x² + 2x + 4 = 100 => 2x² + 2x - 96 = 0 Can you continue to solve for x which is velocity of Rick!

OpenStudy (anonymous):

Uh... I dont really get it...

OpenStudy (anonymous):

I mean: -> 2 x² + 4x + 4 = 100 => 2x² + 4x - 96 = 0

OpenStudy (anonymous):

will it be (2x-12)(x+8)?